OpenStudy (anonymous):

A 45 V battery is connected to a 10.0 Ω resistor and a 5.0 Ω resistor in series. What will be the total current in the circuit? A. 0.33 A B. 3.0 A C. 4.5 A D. 9.0 A

OpenStudy (michele_laino):

the total resistance of your circuit is: R= 10 + 5= ...ohms

OpenStudy (anonymous):

ohh 15?

OpenStudy (michele_laino):

ok! So the current is: I = 45/15=...amperes

OpenStudy (anonymous):

oh! 3? solution is choice B?

OpenStudy (michele_laino):

yes!

OpenStudy (michele_laino):

I = 3.0 Amperes, correct!

OpenStudy (anonymous):

yay! thanks!!:D

OpenStudy (michele_laino):

:)